(3xsquared -2x-1)(xsquared +x +5) multipy and simplify
Answer by lenny460(1073)

(3xsquared -2x -1)(xsquared + x + 5). Multiply and simplify.
The proper way to write is:
(3x^2 - 2x - 1)(x^2 + x + 5)
We then multiply using the FOIL Method:

3x^2 * x^2 = 3x^4
3x^2 * x = 3x^3
3x^2 * 5 = 15x^2
-2x * x^2 = -2x^3
-2x * x = -2x^2
-2x * 5 = -10x
-1 * x^2 = -1x^2
-1 * x = -1x
-1 * 5 = -5

We now have:

3x^4 + 3x^3 + 15x^2 - 2x^3 - 2x^2 - 10x - 1x^2 - 1x - 5


Combine like terms:
3x^3 - 2x^3 = x^3
15x^2 - 2x^2 - 1x^2 = 12x^2
-10x - 1x = -11x


Answer:
3x^4 + x^3 + 12x^2 - 11x - 5









Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com